how do i solve this math problem y=x+5
x+y=11
1
Expert's answer
2011-10-06T08:55:31-0400
This is the system of two linear equations. You can solve it using the
substitution method (for example):
Substitude the expression for y from the
first equation into the second:
x+x+5=11
So, 2x = 6 and x = 3. Then y = x
+ 5 = 3 + 5 = 8.
Answer: x=3, y=8.